A charge of 12 uC when placed in electric field experiences a force of 648 N. What is the magnitude of the electric field strength.

1
Expert's answer
2022-02-13T15:14:38-0500

Given:

"q=12\:\rm \mu C=12*10^{-6}\: C"

"F=648\:\rm N"


The magnitude of the electric field strength

"E=\frac{F}{q}=\frac{648\:\rm N}{12*10^{-6}\:\rm C}=5.4*10^7\rm \: N/C"